OpenStudy (anonymous):

0 to the fifth power is zero right?

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

0^n = 0

myininaya (myininaya):

0^0 is indeterminate form

OpenStudy (dumbcow):

when n is not 0, i should say :)

OpenStudy (anonymous):

thanks

myininaya (myininaya):

:)

OpenStudy (anonymous):

lots of interesting historical debate on \[0^0\]

OpenStudy (anonymous):

thats why i was confused

OpenStudy (anonymous):

0^5 ¨= 0*0*0*0*0 = 0

OpenStudy (anonymous):

interesting being a relative term i guess. http://mathforum.org/dr.math/faq/faq.0.to.0.power.html

OpenStudy (anonymous):

im very new to algebra \[\sqrt[3]{?-125}\] so thats not a real number?

OpenStudy (anonymous):

yes don't get cube roots confused with square roots

OpenStudy (anonymous):

\[\sqrt[3]{-125}=-5\] because \[(-5)^3=-125\]

OpenStudy (anonymous):

you can take the cube root of a negative number. you just cannot take a square root, fourth root, or any even root of a negative number and get a real number back

OpenStudy (anonymous):

The two interesting cases of 0^n are: \[0^0;0^{\infty}\] Both of those are indeterminate. But as long as n is non zero and finite then 0^n=0.

OpenStudy (anonymous):

But why are they interesting?